April 1, 2020, Revision: 1.1

1 Cas A

Figure 1: Top: Q and U, Fourier filtered to k = 80. Bottom: E and B, Fourier filtered to the same k.

$Header: /var/cvs/brandenb/tex/gal/CasA/notes.tex,v 1.1 2020/04/01 17:07:16 brandenb Exp $